If three times the smaller of two consecutive integers is added to four times the larger, the result is 123. Find the smaller integer.
Answer by vidhyak(98)

Let the two consecutive integers be n, n+1.

three times the smaller integer = 3n
four times the larger integer = 4(n+1)

3n + 4(n+1) = 123
3n + 4n + 4 = 123
7n + 4 = 123
7n = 123 - 4 = 119
n = 119/7 =17

smaller integer is 17